OpenStudy (anonymous):

simplify 1-(cos^2x/1+sinx)

OpenStudy (thomas5267):

Is this the thing you need to simplify? \[ 1-\frac{\cos^2(2x)}{1-\sin(x)} \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it is 1+sinx

OpenStudy (anonymous):

sorry it is \[1-\frac{ \cos ^{2}x }{ 1+sinx }\]

OpenStudy (thomas5267):

Hint: \[\cos^2(x)=1-\sin^2(x)\] Then manipulate the expression.

OpenStudy (anonymous):

thank you :)